OpenStudy (accidentalaichan):

Which of the following is not an equivalent form of the compound inequality x+15≤12 and x+15≥5 A. 2 greater than or equal to x + 15 greater than or equal to 5 B. -3 greater than or equal to x greater than or equal to -10 C. A number line with a closed circle on -10, a closed circle on -3, and shading in between. D. A number line with a closed circle on -10, shading to the left, and a closed circle on -3, shading to the right.

OpenStudy (anonymous):

\[x+15\leq 12\\ x\leq -3\]

OpenStudy (anonymous):

\[x+15\geq 5\\ x\geq -10\]

OpenStudy (anonymous):

so \[-10\leq x\leq -3\] shade the stuff between \(-10\) and \(-3\) and use closed circles because of the \(\leq\) instead of \(<\)

OpenStudy (accidentalaichan):

Oh, I see :3